Using the data in this chart what is the primary attack of this disease in the entire population (across all age groups?

Age Group in years

Total Population

Susceptible susceptible population before primary cases occurred

Primary cases

Secondary cases

2 - 4

300

250

100

50

5 - 9

450

420

204

87

10 - 19

152

84

25

15

0.83592%

8.3592

0.436638%

43.634%

83.592%
Answer #1

The correct option is (e): 83.592%

Given that,

The total population of all age groups = 300+450+152= 902

Primary attack of a disease means that chances of the disease to be happen in any person, so susceptible population before the primary stage = 250+420+84= 754

So the percentage= (754 x 100)/902

=83.59%

Thanks.